# A Comprehensive Study of Shapley Value in Data Analytics [Experiment, Analysis & Benchmark]

### v6

@@ 3.2.1. Computation Efficiency
Based on Definition 2 and Eq. 1, the total cost of SV computation is determined by $N_{uc}\times T_{uc}$, where $N_{uc}$ is the total number of utility computations, and $T_{uc}$ is the average time cost of every utility computation. To compute exact SV, the algorithm generally enumerates all possible coalitions of players in the cooperative game by iterations, with each iteration computing the utility on one coalition.
Exact SV computing algorithm is known costly (Rozemberczki et al., 2022; Chen et al., 2023; Bertossi et al., 2023), with the number of iterations reaching an exponential level, i.e., $N_{uc}=2^{n}$. Thus, the computation complexity of exact SV is $\textit{O}(2^{n})$. In DA tasks, the quantity of players (i.e.,$n$) is much higher than expected, resulting in prohibitive computation cost. Moreover, as highlighted by Jethani et al. (2021), the total cost of SV computing in DA can be further increased when DA tasks involve ML models, since each iteration typically necessitates retraining these models, associated with non-trivial expenses. For example, the DO task in Figure 1 analyzes $224\times 224$ players (pixels), thus $N_{uc}=2^{224\times 224}$. Depending on the task’s model size and computation resources, $T_{uc}$ can reach hours to months (Cottier et al., 2024), much longer than milliseconds in traditional game theory.
@@
Table 2. Summary of iteration reduction techniques disentangled from task-agnostic SV computing algorithms. Based on learning paradigms, our work further classifies the DM tasks having applied SV into five types: semi-supervised learning (SSL), active learning (AL), continuous learning (CL), federated learning (FL), and ensemble learning (EL).
@@
Monte Carlo(MC) method computes SV based on Eq. 2 by sampling random permutations iteratively. In each iteration, the marginal contribution of every player $p_{i}(\in\mathcal{N})$ within a randomly sampled permutation, $U(P(\mathcal{O},p_{i}))$, is computed. The final estimation of SV for each player is the average of the marginal contributions calculated for that player in all iterations.
@@
Discussion. Among the sampling-based iteration techniques, MC gains the widest application. Theoretically, GT and CP can further reduce the complexity of SV computation, but these two techniques do not gain much wider application than MC. We speculate that GT is limited by its assumption on the correctness of mirroring the difference between the SV of any two players, while CP is limited by its strong reliance on the sparse SV assumption. The other sampling-based iteration techniques, RE and MLE, are also relatively limited in practical applications compared with MC. These two techniques are heuristic and the corresponding big-O notations are not given by existing arts.
@@
Discussion. Many works (Zhang et al., 2023b; Luo et al., 2024; Azcoitia et al., 2022; Tian et al., 2022a; Cohen et al., 2005; Ghorbani and Zou, 2019; Xia et al., 2024; Pandl et al., 2021; Yang et al., 2022a; Liu et al., 2022a) have attempted the combination of MC+TC to develop hybrid [^1] In this paper, we call the SV computing algorithm adopting different techniques for efficient computation as the hybrid algorithm. SV computing algorithms. As a flexible technique, TC is also compatible with other sampling-based iteration reduction techniques, e.g., RE, MLE, GT, and CP. However, these combinations have not yet been evaluated, since prior works analyze SV applications/computing algorithms as monolithic units rather than decomposing them into reusable building blocks. Later in §5.1, we will evaluate the performance of TC integrated with different sampling-based techniques to validate its flexibility.
@@
Finding 6. For task-specific iteration reduction, leveraging simple-structured ML models or loss-bounded learning algorithms is the key to lowering the complexity of SV computation. However, as the rising popularity of large-scale pre-trained models in the DA domain, the usability of this solution direction might be compromised.
Discussion. Although some task-specific techniques are designed based on DNN, the complexities of those techniques are determined by the number of hidden layers in the model and the number of parameters in each layer. Therefore, if a DA task relies on large-scale pre-trained models (such as GPT-4 (OpenAI, 2023)), those techniques may not work efficiently to generate accurate SV approximation results in this case.
Table 4 summarizes the ML speedup techniques: (1) gradient approximation, (2) test sample skip, and (3) model appraiser. All these techniques are disentangled from task-specific algorithms and are heuristic (thus not given with big-O notations).
@@
Discussion. There are many works on hybrid SV computing algorithms integrating these two types of techniques. An example is MC+GA, utilized by both DV tasks selecting high-quality data tuples from the UK Biobank dataset to train logistic regression (Ghorbani and Zou, 2019) and FL tasks selecting high-quality local models to generate the global model for image classification (Yang et al., 2022a; Liu et al., 2022a; Batool et al., 2022). MC + GA + TSS is another typical combination designed for tasks with the players being ML models (Zheng et al., 2022a). We note that integrating ML speedup with iteration reduction is well-suited for tasks, e.g., DV, FL, where computing a utility involves the costly multi-step gradient descent for model training, or the utility computation relies on a large number of test data samples or complicated models, e.g., pre-trained models with billions of parameters.
@@ 3.2.2. Approximation Error
SV approximate computation, though faster than exact computation, introduces variance due to randomness in sampling the player coalitions and bias caused by incomplete exploration of all the possible coalitions. Hence, it poses a new challenge that the approximate SV may not be an accurate and unbiased estimation of the exact SV, and thus may fail to serve its expected application purposes.
The key to reducing such SV approximation errors is variance reduction. Table 5 summarizes the techniques falling into three types: stratified, antithetic, and kernel-based.
Stratified techniques divide the player coalitions or permutations into several disjoint strata and then randomly select coalitions or permutations from each stratum in a fixed proportion for SV approximation (Wu et al., 2023b; Jia et al., 2019; Zhang et al., 2023a; Watson et al., 2022; Sun et al., 2024; Kolpaczki et al., 2024a; Lin et al., 2023; Otmani et al., 2024; Mitchell et al., 2022; Kolpaczki et al., 2024a, b). The coalitions are stratified based on the number of players included in each coalition, while the permutations are stratified according to the position of a player.
Antithetic techniques generate negatively correlated coalitions or permutations for SV approximation (Otmani et al., 2024; Covert and Lee, 2020; Okhrati and Lipani, 2020; Mitchell et al., 2022). Given a coalition $\mathcal{S}$, its negatively correlated counterpart is $\mathcal{N}-\mathcal{S}$. Given any two permutations, they are negatively correlated if the order of players is completely reversed.
Kernel-based techniques define the distance between permutations by kernels and sample permutations with good distributions relative to kernels when approximating SV (Mitchell et al., 2022).
@@
Discussion. The stratified techniques are applicable regardless of the sampling object, because there exists suitable criteria to divide the coalitions or permutations into disjoint strata. The antithetic techniques also work for sampling the two objects, since given any coalition or permutation, there exists a negatively correlated counterpart. The kernel-based techniques, in contrast, only define the distance and similarity between permutations, and thus cannot work when the sampling object is a coalition.
@@
Discussion. The mainstream algorithms for computing approximate SV rely on sampling. However, according to the law of large numbers (Wikipedia, [n.d.]), no matter which sampling strategy is adopted, it is inevitable to sample more coalitions and compute their utilities for reducing approximation error, which, on the other hand, increases computation complexity. This leaves a riddle: Given a sampling strategy, how to strike a balance between SV approximation error and computation efficiency? Similarly, we conduct certain experiments in §5.2 to bridge this gap.
@@ 3.2.3. Privacy Preservation
Discussion. Lightweight and rigorous theoretical privacy guarantees are two primary factors affecting the selection of privacy protection measures for SV. For applications related to cross-institutional highly-sensitive data (e.g., cross-hospital medical images in Figure 1), rigorous techniques like HE or SMPC are needed. While if computing resources are constrained (e.g., in applications deployed across edge devices), the lightweight NPM, QT, or DR technique is preferred. We note that DP, possessing the lightweight property and rigorous theoretical guarantees simultaneously, is applicable to both cases.
@@
Privacy protection can also influence the effectiveness of SV. The mainstream arts rely on the scaling (Wu et al., 2023a; Salarbashishahri et al., 2022; Sun et al., 2023; Tang et al., 2021; Yang et al., 2023a; Franses et al., 2024; Drungilas et al., 2023) or ranking (Jia et al., 2019; Meepaganithage et al., 2024; Kwon et al., 2021; Ghorbani and Zou, 2019; Zhu et al., 2023a; Tian et al., 2022b) of final SV results to perform pricing, selection, weighting, and attribution in DA. However, the scaled SV results or the ranking results can be easily altered (Bozorgpanah et al., 2022) when integrated with techniques like DP, QT, and DR. We notice a scarcity of quantitative results for answering the question: Can a balance be achieved between the effectiveness of privacy protection with the effectiveness of SV? Therefore, we offer insights about this question with the experimental study in §5.3.
@@ 3.2.4. Interpretability
In addition to the aforementioned challenges, applying SV to DA also faces the trouble of how to properly translate the meaning of the obtained SVs to exact actions (e.g., adding or deleting a data sample, normalizing data features in some dimensions, etc.) in the DA workflow (Kumar et al., 2020).
@@
Counterfactual Given a dataset, there exists a minimum set of tuples in this dataset such that transferring the found set from this dataset to another dataset can flip the direction of the inequality between the SV of those two datasets (Si and Pei, 2024).
@@
Characteristic-based paradigm aims to reveal the relationship between the intrinsic characteristics of each player and its SV. It has been shown that the data tuples with low SV usually belong to noisy data with false labels (Jia et al., 2019), outliers (Ghorbani and Zou, 2019), or corruptions (Ghorbani and Zou, 2019), while the ML models with high SV are those that have high test accuracy and certainty (Xu et al., 2024; Rozemberczki and Sarkar, 2021).
Counterfactual paradigm aims to find out the minimum change between two players that can flip the direction of the inequality between the SVs of those players. Si and Pei (2024) utilized this paradigm to explain why a dataset has a larger SV than another. Given a dataset (defined as the player), the work searched for the minimum set of tuples in the given dataset such that transferring the found set from the given dataset to another dataset can flip the direction of the inequality between the SV of those two datasets.
@@
Discussion. With requirements on expert knowledge and extra computations, the characteristic-based paradigm and the counterfactual explanation did not gain a wider application than the utility-based paradigm. Despite the popularity of utility-based interpretations, many works (Kumar et al., 2020, 2021a; Zhao et al., 2024; Ma and Tourani, 2020) have claimed that the player with a larger SV may have less influence on the overall utility of the associated DA task, which contradicts the mainstream interpretations. However, these works studied only the cases defining data features as the player and thus cannot fully answer the following questions: Can the SV of the four types of players in DA be correctly interpreted by the mainstream paradigm? If cannot, why? Are there general reasons applicable to all four types of players? We endeavor to answer these questions through a comprehensive evaluation in §5.4.
@@ 3.2.5. Summary of Findings
Firstly, our findings, summarized from a wide range of arts, provide general and comprehensive guidelines for academicians and engineers to study and develop SV applications when confronted with DA tasks. Secondly, our findings analyze the arts in this field through a finer-grained perspective. We analyze the pros and cons of the resolution techniques disentangled from complete algorithms in addressing four challenges of using SV in DA (Findings 4, 6, 10, 12), propose the possible combinations of the techniques for developing new algorithms (Findings 5, 7, 8), and discuss the potential conflicts between different challenges, e.g., computation efficiency vs. approximation error (Finding 9) or privacy preservation (Finding 11). Finally, as a consequence of the finer-grained understanding of SV, our findings can instruct the design of a modular and extensible framework for developing SV applications, presented in the next section.
@@ 4. SVBench
Usage Instructions. Users of SVBench, including seasoned engineers and researchers seeking updates on the building blocks of SV applications and newcomers to this filed, need to initialize the framework in three steps. The first step is cooperative game modeling, in which the users properly define the players and the utility function of their targeted DA task. SVBench suggests users perform this step according to Figure 2 and also allows new definitions of player and utility not included in this figure. The second step is SV computation, in which users specify the configuration parameters listed in Table 8. The default settings are marked in the table. The final step is data analysis application/evaluation, in which the users utilize SV results to perform pricing, selection, weighting, or attribution on data or data derivatives according to the specific DA task requirements.
Use Case. Take the DV task in Figure 1 as an example. Suppose the user plans to use SVBench to implement a hybrid SV algorithm, which combines MC and TC techniques with DP for privacy protection, to generate SV of 6000 images from a ‘MedicalImg’ dataset in DV. The user defines 6000 images as the players and a ‘DV-MI’ function, which takes images as inputs to train a classification model and outputs the test accuracy of this model, as the utility function. Then, the user configures SVBench with the following parameters: task=‘DV’, dataset=‘medicalImg’, player=‘tuple’, utility_function=‘DV-MI’, base_algo=‘MC’, optimization_strategy= ‘TC’, privacy_protection_measure=‘DP’, and invokes the SV computing function in SVBench to obtain SV. After that, the user selects the images assigned with the top-50% SV results to learn the classification model. More use cases can be found at GitHub 2.
Extension Supports. Besides implementing the mainstream techniques for computing SV, we provide several APIs in the modules of SVBench (marked by an API icon in Figure 3) for users to extend this framework. One can configure the module he expects to extend by a user-specific parameter at the SV computation step and submit the new functions corresponding to that module. For example, in the above use case, the user submits a user-specific utility function namely DV-MI and configures the utility calculator module by setting utility_function=‘DV-MI’. Similarly, the user can also configure a user-specific aggregator module by submitting a new privacy protection function (e.g., namely newMaskSV, which takes original SV computing results as inputs and outputs masked SV results) and setting privacy_protection_measure=‘newMaskSV’. SVBench will check the legitimacy and validity of the received functions and use the valid functions to execute the operations in the corresponding modules. Moreover, with user permissions, SVBench will embed the valid new functions into their corresponding modules to provide more development choices for future use.
Summary. The application of SV in DA requires to be flexible in engineering implementation, yet most existing works overlook such demands. The application should feature a modular architecture with configurable parameters, empowering engineers to tailor the usage of SV according to domain-specific requirements. Moreover, the application should possess the ability to be parallelized and disaggregated. In this manner, it can leverage diverse computing architectures, including multi-core CPUs, GPUs, and distributed computing paradigms, to enhance scalability and efficiency. With these considerations, we propose SVBench. While SVBench has demonstrated usability, modularity, and flexibility through the success of implementing dozens of algorithms in the next section, we anticipate future extensions to further enhance its potential to develop efficient, secure, and effective SV applications in DA.
5. Evaluation
@@
Base_Algo Sampling Optimization Privacy
@@
RI Iris (Fisher, 1936) Data Feature ($n=4$) Model Output
Wine (Markelle Kelly, [n.d.]) Data Feature ($n=13$)
Adult (Markelle Kelly, [n.d.]) Data Feature ($n=14$)
Ttt (Markelle Kelly, [n.d.]) Data Feature ($n=9$)
@@
DV Iris (Fisher, 1936) Data Tuple ($n=18$) Test Accuracy
Wine (Markelle Kelly, [n.d.]) Data Tuple ($n=15$)
Bank (Markelle Kelly, [n.d.]) Data Tuple ($n=18$)
Ttt (Markelle Kelly, [n.d.]) Data Tuple ($n=22$)
@@
Bank (Markelle Kelly, [n.d.])
Dota2 (Markelle Kelly, [n.d.])
@@
Adult (Markelle Kelly, [n.d.])
Dota2 (Markelle Kelly, [n.d.])
@@
Table 10. Task settings. $n$: total number of players. Each dataset is split into a training set with 80% data samples from the original dataset and a test set with the remaining 20% samples.

### v7

@@ 3.2.1. Computation Efficiency
Based on Definition 2 and Equation 1, the total cost of SV computation is determined by $N_{uc}\times T_{uc}$, where $N_{uc}$ is the total number of utility computations, and $T_{uc}$ is the average time cost of every utility computation. To compute exact SV, the algorithm generally enumerates all possible coalitions of players in the cooperative game by iterations, with each iteration computing the utility on one coalition.
Exact SV computing algorithm is known costly (Rozemberczki et al., 2022; Chen et al., 2023; Bertossi et al., 2023), with the number of iterations reaching an exponential level, i.e., $N_{uc}=2^{n}$. Thus, the computation complexity of exact SV is $\textit{O}(2^{n})$. In DA tasks, the quantity of players (i.e.,$n$) is much higher than expected, resulting in the prohibitive computation cost. Moreover, as highlighted by Jethani et al. (2021), the total cost of SV computing in DA can be further increased when DA tasks involve ML models, since each iteration typically necessitates retraining these models, associated with non-trivial expenses. For example, the DO task in Figure 1 analyzes $224\times 224$ players (pixels), thus $N_{uc}=2^{224\times 224}$. Depending on the task’s model size and computation resources, $T_{uc}$ can reach hours to months (Cottier et al., 2024), much longer than milliseconds in traditional game theory.
@@
Table 2. Summary of iteration reduction techniques disentangled from task-agnostic SV computing algorithms. Based on learning paradigms, our work further classifies the DM tasks having applied SV into five types: semi-supervised learning (SSL), active learning (AL), continuous learning (CL), ensemble learning (EL), and federated learning (FL).
@@
Monte Carlo(MC) method computes SV based on Equation 2 by sampling random permutations iteratively. In each iteration, the marginal contribution of every player $p_{i}(\in\mathcal{N})$ within a randomly sampled permutation, $U(P(\mathcal{O},p_{i}))$, is computed. The final estimation of SV for each player is the average of the marginal contributions calculated for that player in all iterations.
@@
Discussion. Among the sampling-based iteration techniques, MC gains the widest application. Theoretically, GT and CP can further reduce the complexity of SV computation, but these two techniques do not gain much wider application than MC. We speculate that GT is limited by its assumption on the correctness of mirroring the difference between the SVs of any two players, while CP is limited by its strong reliance on the sparse SV assumption. The other sampling-based iteration techniques, RE and MLE, are also relatively limited in practical applications compared with MC. These two techniques are heuristic and the corresponding big-O notations are not given by existing arts.
@@
Discussion. Many works (Zhang et al., 2023b; Luo et al., 2024; Azcoitia et al., 2022; Tian et al., 2022a; Cohen et al., 2005; Ghorbani and Zou, 2019; Xia et al., 2024; Pandl et al., 2021; Yang et al., 2022a; Liu et al., 2022a) have attempted the combination of MC+TC to develop hybrid [^1] In this paper, we call the SV computing algorithm adopting different techniques for efficient computation as the hybrid algorithm. SV computing algorithms. As a flexible technique, TC is also compatible with other sampling-based iteration reduction techniques, e.g., RE, MLE, GT, and CP. However, these combinations have not been evaluated yet, since prior works analyze SV applications or computing algorithms as monolithic units rather than decomposing them into reusable building blocks. Later in §5.1, we will evaluate the performance of TC integrated with different sampling-based techniques to validate its flexibility.
@@
Finding 6. For task-specific iteration reduction, leveraging simple-structured ML models or loss-bounded learning algorithms is the key to lowering the complexity of SV computation. However, with the rising popularity of large-scale pre-trained models in the DA domain, the usability of this solution direction might be compromised.
Discussion. Although some task-specific techniques are designed based on DNN, the complexity of those techniques is determined by the number of hidden layers in the model and the number of parameters in each layer. Therefore, if a DA task relies on large-scale pre-trained models (such as GPT-4 (OpenAI, 2023)), those techniques may not work efficiently to generate accurate SV approximation results in this case.
Table 4 summarizes the ML speedup techniques: gradient approximation, test sample skip, and model appraiser. All these techniques are disentangled from task-specific algorithms and are heuristic (thus not given with big-O notations).
@@
Discussion. There are many works on hybrid SV computing algorithms integrating these two types of techniques. An example is MC+GA, utilized by both DV tasks selecting high-quality data tuples from the UK Biobank dataset to train logistic regression (Ghorbani and Zou, 2019) and FL tasks selecting high-quality local models to generate the global model for image classification (Yang et al., 2022a; Liu et al., 2022a; Batool et al., 2022). MC + GA + TSS is another typical combination designed for tasks with the players being ML models (Zheng et al., 2022a). We note that integrating ML speedup with iteration reduction is well-suited for tasks, like DV or FL, which have (one of) the following characteristics. The first characteristic is that computing a utility involves the costly multi-step gradient descent for model training. Another characteristic is that the utility computation relies on numerous test data samples or complicated models, e.g., pre-trained models with billions of parameters.
@@ 3.2.2. Approximation Error
SV approximate computation, though faster than exact computation, introduces the variance caused by the randomness in sampling the player coalitions and the bias caused by incomplete exploration of all the possible coalitions. Hence, it poses a new challenge that the approximate SV may not be an accurate and unbiased estimation of the exact SV, and thus may fail to serve its expected application purposes. The key to reducing the SV approximation error is variance reduction. Table 5 summarizes the techniques falling into three types: stratified, antithetic, and kernel-based.
Stratified technique divides the player coalitions or permutations into several disjoint strata and then randomly selects coalitions or permutations from each stratum in a fixed proportion for SV approximation (Wu et al., 2023b; Jia et al., 2019; Zhang et al., 2023a; Watson et al., 2022; Sun et al., 2024; Kolpaczki et al., 2024a; Lin et al., 2023; Otmani et al., 2024; Mitchell et al., 2022; Kolpaczki et al., 2024a, b). The coalitions are stratified based on the number of players included in each coalition, while the permutations are stratified according to the position of a player.
Antithetic technique generates negatively correlated coalitions or permutations for SV approximation (Otmani et al., 2024; Covert and Lee, 2020; Okhrati and Lipani, 2020; Mitchell et al., 2022). Given a coalition $\mathcal{S}$, its negatively correlated counterpart is $\𝒮\mathcal{N}\textbackslash\mathcal{S}$. Given any two permutations, they are negatively correlated if the order of players is completely reversed.
Kernel-based technique defines the distance between permutations by kernels and sample permutations with good distributions relative to kernels when approximating SV (Mitchell et al., 2022).
@@
Discussion. The stratified technique is applicable regardless of the sampling objects, because there exist suitable criteria to divide coalitions or permutations into disjoint strata. The antithetic technique also works for sampling the two objects, since given any coalition or permutation, there exists a negatively correlated counterpart. The kernel-based technique, in contrast, only defines the distance and similarity between permutations, and thus cannot work when the sampling object is a coalition.
@@
Discussion. The mainstream algorithms for computing approximate SV rely on sampling. However, according to the law of large numbers (Wikipedia, 2025), no matter which sampling strategy is adopted, it is inevitable to sample more coalitions and compute their utilities to reduce approximation error, which, on the other hand, increases computation complexity. This leaves a riddle: Given a sampling strategy, how to strike a balance between SV approximation error and computation efficiency? Similarly, we conduct certain experiments in §5.2 to bridge this gap.
@@ 3.2.3. Privacy Preservation
Discussion. Lightweight and rigorous theoretical privacy guarantees are two primary factors affecting the selection of privacy protection measures for SV. For applications related to cross-institutional highly-sensitive data (e.g., cross-hospital medical images in Figure 1), rigorous techniques like HE or SMPC are needed. While if computing resources are constrained (e.g., in applications deployed across edge devices), the lightweight NPM, QT, or DR techniques are preferred. We note that DP, possessing the lightweight property and rigorous theoretical guarantees simultaneously, is applicable to both cases.
@@
Privacy protection can also influence the effectiveness of SV. The mainstream arts rely on the scaling (Wu et al., 2023a; Salarbashishahri et al., 2022; Sun et al., 2023; Tang et al., 2021; Yang et al., 2023a; Franses et al., 2024; Drungilas et al., 2023) or ranking (Jia et al., 2019; Meepaganithage et al., 2024; Kwon et al., 2021; Ghorbani and Zou, 2019; Zhu et al., 2023a; Tian et al., 2022b) of final SV results to perform pricing, selection, weighting, and attribution in DA. However, the scaled SV results or the ranking results can be easily altered (Bozorgpanah et al., 2022) when integrated with techniques like DP, QT, and DR. We notice a scarcity of quantitative results for answering the question: Can a balance be achieved between the effectiveness of privacy protection with the effectiveness of SV? Therefore, we offer insights into this question with the experimental study in §5.3.
@@ 3.2.4. Interpretability
In addition to the aforementioned challenges, applying SV to DA also faces the trouble of how to properly translate the obtained SVs to exact actions (e.g., adding or deleting a data sample, normalizing data features in some dimensions, etc.) in the DA workflow (Kumar et al., 2020).
Characteristic-based paradigm seeks to reveal the relationship between the intrinsic characteristics of each player and its SV. It has been shown that the data tuples with low SV usually belong to noisy data with false labels (Jia et al., 2019), outliers (Ghorbani and Zou, 2019), or corruptions (Ghorbani and Zou, 2019), while the ML models with high SV are those that have high test accuracy and certainty (Xu et al., 2024; Rozemberczki and Sarkar, 2021).
Counterfactual paradigm aims to find out the minimum change between two players that can flip the direction of the inequality between the SVs of those players. Si and Pei (2024) utilized this paradigm to explain why a dataset has a larger SV than another. Given a dataset (defined as the player), the work searched for the minimum set of tuples in the given dataset such that transferring the found set from the given dataset to another dataset can flip the direction of the inequality between the SVs of those two datasets.
@@
Counterfactual Given a dataset, there exists a minimum set of tuples in this dataset such that transferring the found set from this dataset to another dataset can flip the direction of the inequality between the SVs of those two datasets (Si and Pei, 2024).
@@
Discussion. With requirements on expert knowledge or extra computations, the characteristic-based paradigm and the counterfactual explanation did not gain a wider application than the utility-based paradigm. Despite the popularity of utility-based interpretations, many works (Kumar et al., 2020, 2021a; Zhao et al., 2024; Ma and Tourani, 2020) have claimed that the player with a larger SV may have less influence on the overall utility of the associated DA task, which contradicts the mainstream interpretations. However, these works studied only the cases defining data features as the player and thus cannot fully answer the following questions: Can the SVs of the four types of players in DA be correctly interpreted by the mainstream paradigm? If cannot, why? Is there any general reason applicable to all four types of players? We endeavor to answer these questions through a comprehensive evaluation in §5.4.
@@ 3.2.5. Summary of Findings
Firstly, our findings, summarized from a wide range of the arts, provide general and comprehensive guidelines for academicians and engineers to study and develop SV applications when confronted with DA tasks. Secondly, our findings analyze the arts in this field through a finer-grained perspective. We analyze the pros and cons of the resolution techniques disentangled from complete algorithms in addressing four challenges of using SV in DA (Findings 4, 6, 10, 12), propose the possible combinations of the techniques for developing new algorithms (Findings 5, 7, 8), and discuss the potential conflicts between different challenges, e.g., computation efficiency vs. approximation error (Finding 9) or privacy preservation (Finding 11). Finally, as a consequence of the finer-grained understanding of SV, our findings can instruct the design of a modular and extensible framework for developing SV applications, presented in the next section.
@@ 4. SVBench
Usage Instructions. Users of SVBench, including seasoned engineers and researchers seeking updates on the building blocks of SV applications and newcomers to this field, need to initialize the framework in three steps. The first step is cooperative game modeling, in which the users properly define the players and the utility function of their targeted DA task. SVBench suggests users perform this step according to Figure 2 and also allows new definitions of player and utility not included in this figure. The second step is SV computation, in which users specify the configuration parameters listed in Table 8. The default settings are marked in the table. The final step is data analysis application/evaluation, in which users utilize SV results to perform pricing, selection, weighting, or attribution on data or data derivatives according to the specific DA task requirements.
Use Case. Take the DV task in Figure 1 as an example. Suppose the user plans to use SVBench to implement a hybrid SV algorithm, which combines MC and TC techniques with DP for privacy protection, to generate SVs of 6000 images from a ‘MedicalImg’ dataset in DV. The user defines 6000 images as the players and a ‘DV-MI’ function, which takes images as inputs to train a classification model and outputs the test accuracy of this model, as the utility function. Next, the user configures SVBench with the following parameters: task=‘DV’, dataset=‘medicalImg’, player=‘tuple’, utility_function=‘DV-MI’, base_algo=‘MC’, optimization_strategy= ‘TC’, privacy_protection_measure=‘DP’, then invokes the SV computing function in SVBench to obtain SVs. After that, the user selects the images assigned with the top-50% SV results to learn the classification model. More use cases can be found at GitHub 2.
@@
Base Algo Sampling Optimization Privacy
@@
Extension Supports. Besides implementing the mainstream techniques for computing SV, we provide several APIs in the modules of SVBench (marked by an API icon in Figure 3) for users to extend this framework. One can configure the module he expects to extend by a user-specific parameter at the SV computation step and submit the new functions corresponding to that module. For example, in the above use case, the user submits a user-specific utility function namely DV-MI and configures the utility calculator module by setting utility_function=‘DV-MI’. Similarly, the user can also configure a user-specific aggregator module by submitting a new privacy protection function (e.g., namely newMaskSV, which takes original SV computing results as inputs and outputs masked SV results) and setting privacy_protection_measure=‘newMaskSV’. SVBench will check the legitimacy and validity of the received functions and use the valid functions to execute the operations in the corresponding modules. Moreover, with the user permission, SVBench will embed the valid new functions into their corresponding modules to provide more development choices for future use.
Summary. The application of SV in DA requires to be flexible in engineering implementation, yet most existing works overlook such demands. The application should feature a modular architecture with configurable parameters, empowering engineers to tailor the usage of SV according to task-specific requirements. Moreover, the application should possess the ability to be parallelized and disaggregated. In this manner, it can leverage diverse computing architectures, including multi-core CPUs, GPUs, and distributed computing paradigms, to enhance scalability and efficiency. With these considerations, we propose SVBench. While SVBench has demonstrated usability, modularity, and flexibility through the success of implementing dozens of algorithms in the next section, we anticipate future extensions to further enhance its potential to develop efficient, secure, and effective SV applications in DA.
@@
RI Iris (Markelle Kelly, 2025) Data Feature ($n=4$) Model Output
Wine (Markelle Kelly, 2025) Data Feature ($n=13$)
Adult (Markelle Kelly, 2025) Data Feature ($n=14$)
Ttt (Markelle Kelly, 2025) Data Feature ($n=9$)
@@
DV Iris (Markelle Kelly, 2025) Data Tuple ($n=18$) Test Accuracy
Wine (Markelle Kelly, 2025) Data Tuple ($n=15$)
Bank (Markelle Kelly, 2025) Data Tuple ($n=18$)
Ttt (Markelle Kelly, 2025) Data Tuple ($n=22$)
@@
Bank (Markelle Kelly, 2025)
Dota2 (Markelle Kelly, 2025)
@@
Adult (Markelle Kelly, 2025)
Dota2 (Markelle Kelly, 2025)
@@
Table 10. Task settings. $n$: total number of players. Each dataset is split into a training set with 80% samples from the original dataset and a test set with the remaining 20% samples.
5. Evaluation